OpenStudy (anonymous):

Solve by substitution 7a = 2 - b 7a= 6 - b a.no solution b.infinite solutions c. ordered pair:

OpenStudy (aroub):

7a = 2 - b 7a= 6 - b b=2-7a 7a=6-2+7a 7a-7a=4 0=4 So, no solution.

mathslover (mathslover):

OR \[\large{2-b=6-b \quad \textbf{since {=7a}}}\] \[\large{2-6=-b+b}\] \[\large{-4\ne 0}\] hence no soln

mathslover (mathslover):

@aroub explained well though

mathslover (mathslover):

got it @Jenny4218 ?